OpenStudy (anonymous):

How can the cotangent of an angle in a right triangle be calculated? by dividing the hypotenuse by the adjacent side by dividing the adjacent side by the opposite side by dividing the adjacent side by the hypotenuse by dividing the hypotenuse by the opposite side

OpenStudy (anonymous):

sin=opp/hyp cos=adj/hyp tan=opp/adj cotan=1/tan can you do it now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1/tan= 1/(opp/adj)= adj/opp

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome